What is the temperature of an 11.2 L sample of carbon monoxide, CO, at 744 torr if it occupies 13.3 L at 55°C and 744 torr?

1
Expert's answer
2022-03-12T02:01:09-0500

The pressure is constant, so we can use Charles’ Law:

 V1/T1 = V2/T2

Invert both sides of the equation:

T1/V1 = T2/V2

Multiply both sides by V2:

T2 = T1 x V2/V1

V1 = 13.3 L; T1 = (55 + 273.15) K = 328.15 K

V2 = 11.2 L; T2 = ?

T2 = 328.15 K x 11.2 L/13.3 L = 276 K = 3°C